# Neural Surface Priors for Editable Gaussian Splatting

### v1

@@ 3 Method
$\mathbf{r}_{0}^{\prime}=\frac{(\mathbf{v}_{1}^{\prime}-\mathbf{v}_{0}^{\prime})\times(\mathbf{v}_{2}^{\prime}-\mathbf{v}_{0}^{\prime})}{\|(\mathbf{v}_{1}^{\prime}-\mathbf{v}_{0}^{\prime})\times(\mathbf{v}_{2}^{\prime}-\mathbf{v}_{0}^{\prime})\|}\,,\ \mathbf{r}_{1}^{\prime}=\frac{\mathbf{v}_{1}^{\prime}-\mathbf{v}_{0}^{\prime}}{\|\mathbf{v}_{1}^{\prime}-\mathbf{v}_{0}^{\prime}\|}\,,$
and $\mathbf{r}_{2}^{\prime}$ is the orthonormal projection of $\mathbf{v}_{2}^{\prime}-\mathbf{v}_{0}^{\prime}$ onto the plane spanned by $\mathbf{r}_{0}^{\prime}$ and $\mathbf{r}_{1}^{\prime}$. The corresponding scaling parameters are then $s_{0}=\varepsilon$, $s_{1}=\|\mathbf{v}_{1}^{\prime}-\mathbf{v}_{0}^{\prime}\|$, and $s_{2}=\langle\mathbf{v}_{2}^{\prime}-\mathbf{v}_{0}^{\prime},\mathbf{r}_{2}^{\prime}\rangle$. We refer the reader to [23] for more details.
@@
$\mathbf{u}_{0}=\frac{\mathbf{w}_{1}-\mathbf{w}_{0}}{\|\mathbf{w}_{1}-\mathbf{w}_{0}\|}\,,\ \mathbf{u}_{1}=\frac{(\mathbf{w}_{1}-\mathbf{w}_{0})\times(\mathbf{w}_{2}-\mathbf{w}_{0})}{\left\|(\mathbf{w}_{1}-\mathbf{w}_{0})\times(\mathbf{w}_{2}-\mathbf{w}_{0})\right\|}\,,$
@@
$\mathbf{V}^{\prime}=\mathbf{T}\left(\mathbf{V}-\mathbf{w}_{0}\right)+\mathbf{w}^{\prime}_{0}\,,$
where we first subtract the reference vertex $\mathbf{w}_{0}$ to recenter the triangle vertices, and then translate the modified triangle by $\mathbf{w}^{\prime}_{0}$ to position it in its new location. Note that the exact values of $\mathbf{T}$, $\mathbf{w}$ and $\mathbf{w^{\prime}}$ depend on the triangle $\mathbf{V}\in\mathbf{T}$ being currently processed.
@@
4 Experiments
Figure 5: Example modifications produced with our method for scenes in the NeRF Synthetic dataset. First row: ficus with a repositioned branch. Second row: various rotations applied to objects in the materials scene. Third row: the drums scene with a displaced cymbal. Fourth: broken masts in the ship scene. These scenes have been edited in Blender using a combination of the translate, rotate, knife, bevel, and randomize tools.
@@ 4 Experiments
Figure 6: Progression of a soft body simulation of a Ficus plant. Top row: changes in the object’s geometry. Bottom row: corresponding modifications in appearance obtained using our method.
Numerical results.
We assess the qualitative performance of our method in the tasks of geometry reconstruction and novel-view synthesis. For geometry reconstruction, we utilize the DTU dataset, employing the Chamfer-$L_{1}$ distance metric to measure the surface quality. The DTU dataset provides 2D ground truth images with a resolution of $1600\times 1200$, along with associated camera poses, and features objects with various material properties. Depending on the selected scene, the dataset includes either $49$ or $64$ RGB images. We compare our approach against several established methods: NeuS [24], NeRF [12], VolSDF [26], Relightable 3D Gaussian (ReGS) [5], 3DGSR [11] and COLMAP [20].
Table 1 demonstrates that our method surpasses other state-of-the-art methods in terms of the surface reconstruction quality, as measured by the Chamfer-$L_{1}$ distance. We tested our method in two scenarios: images with and without masks. The values for the reference methods were calculated for images without masks. Regardless of whether we use a DTU dataset with masks or without masks, our method allows us to obtain meshes of very high quality.
Scan ID 24 37 40 55 63 65 69 83 97 105 106 110 114 118 122 Avg
COLMAP 0.81 2.05 0.73 1.22 1.79 1.58 1.02 3.05 1.40 2.05 1.00 1.32 0.49 0.78 1.17 1.36
NeRF 1.90 1.60 1.85 0.58 2.28 1.27 1.47 1.67 2.05 1.07 0.88 2.53 1.06 1.15 0.96 1.49
NeuS 1.00 1.37 0.93 0.43 1.10 0.65 0.57 1.48 1.09 0.83 0.52 1.20 0.35 0.49 0.54 0.84
VolSDF 1.14 1.26 0.81 0.49 1.25 0.70 0.72 1.29 1.18 0.70 0.66 1.08 0.42 0.61 0.55 0.86
ReGS 1.21 1.00 1.10 0.87 1.03 1.51 1.51 1.06 1.63 0.97 1.36 1.21 0.94 1.38 1.26 1.20
3DGSR 0.68 0.84 0.70 0.39 1.16 0.87 0.77 1.48 1.26 0.87 0.69 0.80 0.42 0.64 0.60 0.81
Ours (w/ mask) 0.52 0.67 0.34 0.37 0.93 0.60 0.57 1.22 0.79 0.67 0.49 0.75 0.36 0.42 0.47 0.61
Ours (w/o mask) 0.52 0.75 0.41 0.37 0.90 0.66 0.59 1.37 1.08 0.85 0.46 0.98 0.33 0.49 0.50 0.68
Table 1: We evaluate the accuracy of surface reconstruction, with color-coded cells denoting the top-performing results with respect to Chamfer-$L_{1}$ distance. Our approach is compared to multiple state-of-the-art methods. It demonstrates superior performance over all competitors, achieving the lowest error.
Chair Drums Ficus Hotdog Lego Materials Mic Ship Avg
NeRF 34.17 25.08 30.39 36.82 33.31 30.03 34.78 29.30 31.74
Ins-NGP 35.00 26.02 33.51 37.40 36.39 29.78 36.22 31.10 33.18
Mip-NeRF 35.14 25.48 33.29 37.48 35.70 30.71 36.51 30.41 33.09
3D-GS 35.36 26.15 34.87 37.72 35.78 30.00 35.36 30.80 33.32
NeuS 31.22 24.85 27.38 36.04 34.06 29.59 31.56 26.94 30.20
NeRO 28.74 24.88 28.38 32.13 25.66 24.85 28.64 26.55 27.48
BakedSDF 31.65 20.71 26.33 36.38 32.69 30.48 31.52 27.55 29.66
NeRF2Mesh 34.25 25.04 30.08 35.70 34.90 26.26 32.63 29.47 30.88
3DGSR 34.85 26.08 35.17 36.88 34.90 30.03 36.44 31.48 33.23
Ours 35.32 26.09 34.32 37.77 35.42 29.22 36.62 30.64 33.18
Table 2: Comparison of performance in novel view synthesis. All methods have been categorized as either SDF-based (top) no-SDF (bottom). In several scenes, our method achieves the highest PSNR score.
For novel-view synthesis, we evaluate the PSNR on the NeRF Synthetic dataset, which consists of eight manually designed scenes. Each scene includes $100$ training images with a resolution of $800\times 800$.
Table 2 shows the novel view synthesis evaluation results. We compare our approach with NeRF [12], Instant NGP (Ins-NGP) [13], Mip-NeRF [1], 3DGSR [11], NeuS [24], NeRO [9], BakedSDF [27], and NeRF2Mesh [21].
Among the SDF-based models, our approach achieves the best results for the chair, drums, hotdog, lego and mic scenes. Across all models in the comparison, we achieve the highest PSNR value for the mic and hotdog scene and the highest average PSNR across all scenes.
Editing 3D objects
To illustrate the applicability of our approach, we conducted a series of adjustments on selected scenes, focusing on edits that mirror that of typical everyday environments, such as reorienting a branch of a plant. To modify the geometry, we used Blender, and then propagated these modifications to the appearance models using our method. The results are presented in Figure 5 and showcase a combination of the translate, rotate, knife, bevel, and randomize Blender tools applied to various scenes. Additionally, we showcase in Figure 7 that our method is capable of modifications directly on the triangle soup with arbitrary transforms such as non-linearly warping the ship scene.
Physics
We assessed the performance of our method in dynamic scenarios by running soft body simulations in Blender. These simulations allow objects to exhibit flexible movements. Using our method, we transferred the resulting dynamics from the surface of the mesh to the object’s appearance, as illustrated in Figure 6. While initially the dynamics are captured effectively, as the simulation progresses, an increasing number of artifacts becomes noticeable.
Figure 7: Our framework also allows to apply arbitrary, non-mesh-guided transforms to the recovered appearance by directly modifying the vertices of proxy triangles, similar to GaMeS [23].
5 Conclusion
In this paper, we introduced a 3DGS-based method for geometry and appearance recovery from ground truth image data, which supports mesh-guided scene modification using a pretrained neural surface prior. Our method aims to maintain high visual fidelity in edits while providing flexibility during the editing process improving the workflow of computer graphics artists.
While this approach allows users to easily modify the recovered appearance by adjusting the underlying mesh geometry, it also faces limitations. First, the editing process is highly dependent on the accuracy of the estimated mesh. Scenes with complex, fine-grained geometric details are challenging for surface extracting methods, consequently making precise mesh editing more difficult. Enhancing the resilience to input inaccuracies is a potential area for improvement. Furthermore, while certain mesh editing operations such as translation, scaling, shearing, rotation or other commonly available operations are well supported by our method, a mesh editing operation resulting in a different number of faces is not supported by our method.
Finally, editing operations such as translations may introduce appearance inconsistencies. An example, would be a shadow embedded in the recovered appearance of objects which would remain after a subcomponent of the reconstructed scene has been moved to a different location. In such cases relighting approaches would have to be used to remove such inconsistencies from the scene.

### v2

@@ III-C Mesh-guided Appearance Modification
$\mathbf{r}_{0}^{\prime}=\frac{(\mathbf{v}_{1}^{\prime}-\mathbf{v}_{0}^{\prime})\times(\mathbf{v}_{2}^{\prime}-\mathbf{v}_{0}^{\prime})}{\|(\mathbf{v}_{1}^{\prime}-\mathbf{v}_{0}^{\prime})\times(\mathbf{v}_{2}^{\prime}-\mathbf{v}_{0}^{\prime})\|}\,,\ \mathbf{r}_{1}^{\prime}=\frac{\mathbf{v}_{1}^{\prime}-\mathbf{v}_{0}^{\prime}}{\|\mathbf{v}_{1}^{\prime}-\mathbf{v}_{0}^{\prime}\|}\,,$ (4)
and $\mathbf{r}_{2}^{\prime}$ is the orthonormal projection of $\mathbf{v}_{2}^{\prime}-\mathbf{v}_{0}^{\prime}$ onto the plane spanned by $\mathbf{r}_{0}^{\prime}$ and $\mathbf{r}_{1}^{\prime}$. The corresponding scaling parameters are then $s_{0}=\varepsilon$, $s_{1}=\|\mathbf{v}_{1}^{\prime}-\mathbf{v}_{0}^{\prime}\|$, and $s_{2}=\langle\mathbf{v}_{2}^{\prime}-\mathbf{v}_{0}^{\prime},\mathbf{r}_{2}^{\prime}\rangle$. We refer the reader to [waczynska2024games] for more details.
@@
$\mathbf{u}_{0}=\frac{\mathbf{w}_{1}-\mathbf{w}_{0}}{\|\mathbf{w}_{1}-\mathbf{w}_{0}\|}\,,\ \mathbf{u}_{1}=\frac{(\mathbf{w}_{1}-\mathbf{w}_{0})\times(\mathbf{w}_{2}-\mathbf{w}_{0})}{\left\|(\mathbf{w}_{1}-\mathbf{w}_{0})\times(\mathbf{w}_{2}-\mathbf{w}_{0})\right\|}\,,$ (5)
@@
$\mathbf{V}^{\prime}=\mathbf{T}\left(\mathbf{V}-\mathbf{w}_{0}\right)+\mathbf{w}^{\prime}_{0}\,,$ (6)
where we first subtract the reference vertex $\mathbf{w}_{0}$ to recenter the triangle vertices, and then translate the modified triangle by $\mathbf{w}^{\prime}_{0}$ to position it in its new location. Note that the exact values of $\mathbf{T}$, $\mathbf{w}$ and $\mathbf{w^{\prime}}$ depend on the triangle $\mathbf{V}\in\mathcal{T}$ being currently processed.
@@
IV Experiments
Figure 5: Comparison of rendering outputs for different Gaussian initialization schemes. Random initialization introduces artifacts, whereas sampling initial locations from the extracted mesh surface ensures cleaner and more accurate results.
@@ IV Experiments
Figure 6: Progression of a wind-driven simulation of a Ficus plant. Top row: changes in the object’s geometry. Bottom row: corresponding modifications in appearance obtained using our method.
IV-A Numerical Results
We assess the qualitative performance of our method in the tasks of geometry reconstruction and novel-view synthesis.
For novel-view synthesis, we evaluate the PSNR on the NeRF Synthetic dataset, which consists of eight manually designed scenes. Each scene includes $100$ training images with a resolution of $800\times 800$ and associated camera poses. Table I shows the results of this evaluation. We compare our approach with NeRF [mildenhall2020nerf], Instant NGP (Ins-NGP) [mueller2022instant], Mip-NeRF [barron2021mipnerf], 3DGSR [lyu20243dgsr], NeuS [wang2023neus], NeRO [liu2023nero], BakedSDF [yariv2023bakedsdf], and NeRF2Mesh [tang2022nerf2mesh]. Among SDF-based models, our approach outperforms others in the chair, drums, hotdog, lego, and mic scenes. Across all compared models, we achieve the highest PSNR for the mic and hotdog scenes.
Morever, we investigated the impact of mesh resolution on the recovered appearance. Table II presents the results of the PSNR, SSIM and LPIPS metrics across different mesh resolutions. The analysis indicates that metric values remain consistent regardless of the mesh resolution, demonstrating that render quality is unaffected by mesh granularity.
Chair Drums Ficus Hotdog Lego Materials Mic Ship
NeRF 34.17 25.08 30.39 36.82 33.31 30.03 34.78 29.30
Ins-NGP 35.00 26.02 33.51 37.40 36.39 29.78 36.22 31.10
Mip-NeRF 35.14 25.48 33.29 37.48 35.70 30.71 36.51 30.41
3D-GS 35.36 26.15 34.87 37.72 35.78 30.00 35.36 30.80
NeuS 31.22 24.85 27.38 36.04 34.06 29.59 31.56 26.94
NeRO 28.74 24.88 28.38 32.13 25.66 24.85 28.64 26.55
BakedSDF 31.65 20.71 26.33 36.38 32.69 30.48 31.52 27.55
NeRF2Mesh 34.25 25.04 30.08 35.70 34.90 26.26 32.63 29.47
3DGSR 34.85 26.08 35.17 36.88 34.90 30.03 36.44 31.48
Ours 35.32 26.09 34.32 37.77 35.42 29.22 36.62 30.64
TABLE I: Novel view synthesis performance comparison.
Resolution SSIM PSNR LPIPS Vertices Faces Time [s]
200 0.9801 35.5185 0.01776 60361 120778 15.77
400 0.9801 35.5331 0.01785 255931 512304 15.76
600 0.9802 35.5607 0.01763 593934 1188712 19.82
800 0.9802 35.5506 0.01765 1050086 2101122 22.93
1000 0.9804 35.5663 0.01754 1667066 3334984 29.53
TABLE II: Comparison of metrics, vertex/face count, and edit time across mesh resolutions for the Lego scene.
Chair Drums Ficus Hotdog Lego Materials Mic Ship
Mesh 35.32 26.09 34.32 37.77 35.42 29.22 36.62 30.64
Random 28.07 24.71 27.53 34.54 31.25 24.47 29.48 24.12
TABLE III: PSNR for different initialization schemes.
Subsequently, we examined how mesh resolution impacts editing quality. Figure 4 compares an unmodified high mesh resolution scene with a modified appearance derived by editing a lower, $100$-resolution mesh. Despite the reduced geometric detail, the appearance retains the visual quality; artifacts in lower-quality meshes do not affect rendering quality unless they directly impact the modifications. Additionally, lower-resolution meshes reduce computational cost and speed up editing (cf. Table II). These findings demonstrate that our method enables efficient editing, regardless of mesh quality.
IV-B Editing 3D Objects
To illustrate the applicability of our approach, we conducted experiments on various scenes from the NeRF Syntetic and BlendedMVS datasets. BlendedMVS[yao2020blendedmvs] consists of 17,000 training samples across 113 diverse scenes, including architectures, sculptures and small objects.
We used Blender for mesh modification, and propagated the edits to the appearance using our method. The results, shown in Figure 7, illustrate the application of various Blender tools (translate, rotate, knife, bevel, and randomize) tools applied to across different scenes.
IV-C Physics
We evaluate our method’s performance in a dynamic scenario by simulating wind physics in Blender. This simulation models air movement affecting leaves and branches of the plant. Using our method, we transferred the resulting dynamics from the surface of the mesh to the object’s appearance, as illustrated in Figure 6. This demonstrates how our method consistently adapts the appearance of the represented object to dynamic deformations.
IV-D Ablation Study
Our method initializes Gaussians by sampling 100,000 points from the extracted mesh. To assess the impact of this initialization scheme, we instead initialized with 100,000 uniformly distributed points within a bounding cube. Table III shows the evaluation results, showing that surface-based sampling yields higher PSNR scores. Additionally, visual analysis of the rendered outputs (Fig. 5) reveals that random initialization introduces noticeable artifacts.
Figure 7: Example modifications produced with our method for scenes in the NeRF Synthetic and BlendedMVS datasets.
V Conclusions
In this paper, we introduced a 3DGS-based scene editing method that enables mesh-guided modifications of the recovered appearance. Our approach ensures edits remain visually consistent by leveraging a scene representation trained in two steps: first, a a neural surface is trained to approximate scene geometry, which is then used as a prior for guiding the alignment of Gaussians. The structure allows for flexible and intuitive editing while addressing the limitations of triangle soup representations, enabling a broader scope of topology-aware modifications.
While our method effectively propagates mesh edits to the recovered appearance, it is not without limitations. First, the underlying mesh must provide a structurally sound approximation of the scene geometry. First, the underlying mesh must provide a sufficiently consistent approximation of the scene geometry. While it does not need to be an exact match, issues such as disconnected components or floating artifacts in the geometry can introduce inconsistencies in the editing process, highlighting the importance of the neural surface prior.
Second, our approach assumes a fixed mesh topology. While transformations such as scaling, shearing, translation and rotation are well supported, editing operations that modify the number of mesh faces, such as remeshing or topology refinement, are not integrated into our framework, as we rely on maintaining correspondence between original and edited mesh faces. Future work could explore ways to accommodate dynamic mesh topology.
Finally, certain edits may introduce appearance inconsistencies, particularly in lightning-dependent effects such as embedded shadows. For example, if an object is moved, its original shadow remains unchanged in the modified appearance. Addressing such inconsistencies would require relighting techniques, which are a natural extension of our approach.
[…]